Statistics south Africa reported the fo3 data fir 2012 labour force 17 948 000. Employment, 13 422 000, working age population, 32 786 000. Calculate the unemployment rate, labour force participant rate and absorption rate


1
Expert's answer
2021-09-07T09:46:48-0400

Given here the data for the labour force of 2012:

Labour force: 17,948,000

Employment: 13,422,000

Working Age Population: 32,786,000


Hence, the following can be calculated. 

The unemployment rate can be defined as the part of the labour force that is unemployed but actively looking for a job over the last 30 days. It can be calculated with the following formula:

"Unemployment\space Rate =\frac{ Number \space of\space people\space Unemployed}{

Labour\space Force

}

×100"


Since the number of unemployed people is not given, it can be calculated as the part of the labour force that is not employed, with the help of the following formula:

"Labour \space Force = Employed+Unemployed"


Here, the labour force and the number of employed is given, which will help us calculate the number of unemployed.

"Labour \space Force = Employed+Unemployed\\17,984,000=13,422,000+Unemployed\\Unemployed=17,984,000-13,422,00\\=4,586,000"


Hence, the unemployment rate is:

"Unemployment \space Rate =\frac{ 4,596,000}{17,984,000}×100=25.5\%"

The unemployment rate is 25.5% in the economy. 

 

 The labour-force participation rate is a statistical measure used to indicate an economy's entire working population. This consists of the employed as well as the unemployed, looking for work. Hence, it is the part of the working population engaged in active labour or searching for it. The calculation can be done as:

"Labour\space Force\space Participation = \frac{Labour\space Force}{Workingn\space Age\space Population}×100\\=\frac{17,984,000}{32,786,000}\\=0.55"

The labour force participation rate is 0.55.